OpenStudy (anonymous):

Help pleasee!! How do you change 8/5 into a radical?!?

OpenStudy (jdoe0001):

8/5 into a radical? what is that?

OpenStudy (anonymous):

what is what a radical?

OpenStudy (jdoe0001):

yes, I mean.. what's the expected answer anyhow? how should it look like?

OpenStudy (welshfella):

i was going to ask same question as jdoe0001

OpenStudy (e.mccormick):

A radical or a ratio? Because a radical is like \(x\sqrt{y}\)

OpenStudy (anonymous):



OpenStudy (jdoe0001):

hmmmm ok

OpenStudy (e.mccormick):

Is the original this? \(\dfrac{8}{\sqrt{5}}\)

OpenStudy (anonymous):

no its just 8/5

OpenStudy (e.mccormick):

Or something like that? Perhaps: \(\dfrac{\sqrt{8}}{\sqrt{5}}\) Because \(\dfrac{8}{5}\ne \dfrac{2\sqrt{10}}{5}\)

OpenStudy (anonymous):

no its just 8/5

OpenStudy (e.mccormick):

See, \(2\sqrt{10} \approx 6.3245553203367586639977870888654\) Hey, did they do this? \(\sqrt{\dfrac{8}{5}}\)

OpenStudy (anonymous):

wait what

OpenStudy (jdoe0001):

yeah...... is not quite workable for one \(\bf \cfrac{8}{5}=1.6\qquad \cfrac{2\sqrt{10}}{5}=1.2649110640673517328 \\ \quad \\ 1.6\ne 1.2649110640673517328\)

OpenStudy (jdoe0001):

check your material, maybe something is amiss or post a quick screenshot of the material

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

i cant get it to screen shot but we are doing geometric means. we are given the numbers 4/5 and 2. i have my proportions set up as 4/5/x=x/2 and got 8/5

OpenStudy (jdoe0001):

if it's on the screen... press PrintScr button, usually next to the [F12] button on the keyboard then go to Paint and Edit > paste crop as needed, annotate as needed, save it and post it here

OpenStudy (e.mccormick):

OH! Geometric means! Those put everything under a radical!

OpenStudy (anonymous):

so like can u help me then?

OpenStudy (welshfella):

geometric mean of a and b is sqrt(ab)

OpenStudy (e.mccormick):

Is the actual question this? The geometric mean of 4, 5, and x = the geometric mean of x and 2

OpenStudy (anonymous):

no the question has a pair of numbers. the numbers are 4/5 and 2

OpenStudy (anonymous):

to get a proportion we were suppost to do 4/5/x=x/2

OpenStudy (e.mccormick):

OK! yah, then take 4/5 as a and 2 as b and put it into what welshfella said.

OpenStudy (e.mccormick):

For n constants k, (\(k_1, k_2, \dots ,k_n\)) the geometric mean is: \(\sqrt[n]{k_1\times k_2\times \dots \times k_n}\) So you put them under the radical, multiply, then take the nth root. Because you have two, it is the square root.

OpenStudy (anonymous):

i have no clue how to do that. she never taught us..

OpenStudy (e.mccormick):

Well, that is how it is done. For a geometric mean of two things, you multiply the two things than take theor square root. If it was three seperate things (the fraction is 1 thing, not two) then you would multiply the three things and take their cube root. For 4 things, multiply the 4 then take the 4th root. On and on.

OpenStudy (anonymous):

im just wondering how it went from 8/5 to 2 square root 10 over 5

OpenStudy (e.mccormick):

because it is a geometric mean! A geometric mean turns everything into a radical of some sort.

OpenStudy (e.mccormick):

See if this makes some sense: http://www.freemathhelp.com/geometric-mean.html